OpenStudy (anonymous):

assign oxidation number to underlined atoms so2, s is the underlined and k2cr2o7, cr2 is the underlined

OpenStudy (anonymous):

well you know oxygen can only have a -2 charge, thus 2 oxyens x -2 = -4 on the sulfur atom k2cr2o7, potassium can only be +1 charge, so 2k = +2 charge; oxygen -2, 7x-2 = -14 -14 + 2 + 2(chromium's charge) = 0 solve for cr we get +6